OpenStudy (anonymous):

One more rectangular-shaped piece of metal siding needs to be cut to cover the exterior of a pole barn. The area of the piece is 30 ft². The length is 1 less than 3 times the width. How wide should the metal piece be? Round to the nearest hundredth of a foot.

OpenStudy (anonymous):

3.33 ft 4.3 ft 1 ft 30 ft

OpenStudy (anonymous):

a

OpenStudy (anonymous):

\[\{L w=30,L=3w-1\}\]\[\left\{L=9,w=\frac{10}{3}\right\} \]